answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many ways can 8 books be arranged on a bookshelf?

Assuming they are all right side up and spine out, there would be 40,320 different ways of arranging them.

Consider: There are 8 choices for the 1st book, 7 for the 2nd, 6 for the 3rd, 5 for the 4th, 4 for the 5th, 3 for the 6th, 2 for the 7th and 1 for the 8th. Multiplying these numbers out:

8 x 7 x 6 x 5 x 4 x 3 x 2 x 1 = 40,320. More than you thought?

The mathematical name for this is called finding the factorial of a number!

It is written n! ( n followed by an explanation point) -

Are the ages of people in the united states normally distributed?

No. Ages of people dicrease as the age increases - not just in the US but in all countries.

In a certain distribution the mean is 50 with a standard deviation of 4 tell the probability that a number lies between 42 and 68?

Calculate the Z value and determine the area (probability) between the Z values. Z = (x-mu)/sigma. Z42 = (42-50)/4 = -2 & Z68 = (68-50)/4 = 4.5. Go to the related link and the areas are .0228 & for 4.5 use .9999. Subtract the two values; .9999 - .0228 = .9771.

Why is the Normal Distribution called normal?

Suppose you could call it the Gaussian Distribution or the Laplace-Gauss (not to be confused with the Laplace distribution which takes an absolute difference from the mean rather than a squared error)... however the Brits had no one to name this distribution after (not the German and French names) and because it is the ubiquitous distribution they just called it... well the NORMAL!!

A bag contain 4 green 6 black 7 white balls a ball is drawn at random what is the probability that it is a green or a black ball?

This is a law of addition probability which states that the probability of A or B equals the probability of A plus the probability of B minus the probability of A and B. Written in mathematical terms, the equation is: P(AorB) = P(A) + P(B) - P(AnB) where P(AnB) = 0 (since you can not pull out a green and black ball at the same time). Let P(A) = Probability of drawing the green ball & let P(B) = Probability of drawing the black ball. Total outcomes is 17. So, P(A) = 4/17 & P(B) = 6/17. Therefore P(green or black) = 4/17 + 6/17 = 10/17.

What is the probability of rolling the sum of 6 with 2 dice?

there is a 1 to 5 probablilty that you will roll the sum of 6 with two dice.

If two events are independent the probability that both occur is?

That probability is the product of the probabilities of the two individual events; for example, if event A has a probability of 50% and event B has a probability of 10%, the probability that both events will happen is 50% x 10% = 5%.

Find the probability of rolling a sum of 9 with two dice?

There is 4 ways to roll a 9 with 2 dice, and 36 possible outcomes. So, the probability of rolling a sum of 9 with two dice is 4/36 or 1/9.

Is it possible to have a probability greater than 100?

It is not possible to have a probability greater than 1. All probabilities are between 0 and 1, inclusive.

How many possible outcomes if you roll a die 4 times?

It depends what you mean by outcome. If you are interested in the total number of different outcomes, then there are 6 x 6 x 6 x 6 or 1296 unique possible outcomes. If you are only interested in the total (the sum) of the resulting roll, it can be anything from 4 to 24 so there are just 21 possible outcomes.

Find two consecutive integers whose sum is 19?

Two consecutive integers whose sum is 19 is 10 & 9; 10+9 = 19.

If a multiple choice test has five answers what is the probability of guessing the right answer?

You have a 50% chance of getting the right answer and a 50% of getting the wrong answer. You best bet is to eliminate your three worse/wrong answers and that leaves you with two that are more or less correct . I have taken tests where points are taken away for guessing. Each right answer was one point and for every wrong two points were taken off. This was a national standardized test. In all the testing I have taken I have no way to beat the system set up by testing. I have taken the LSAT, GRE, SAT, ACT, National Teachers Exam, and several others. Nothing beats getting the test prep book and learning the tricks for that particular test.

What are the odds of two people having same birthday AND birthyear?

1 in 20,000!

I know 4 people with my same birthday, 3 of them not being related...

Relationship between rayleigh distribution and gaussian distribution?

If X and Y are independent Gaussian random variables with mean 0 and standard deviation sigma, then sqrt(X^2 + Y^2) has a Rayleigh distribution with parameter sigma.

If a coin is flipped twice what is the probability that both will be heads?

The sample space is HH, HT, TH, HH. Since the HH combination can occur once out of four times, the probability that if a coin is flipped twice the probability that both will be heads is 1/4 or 0.25.